OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = x2 + 1,find f(g(3)).

Directrix (directrix):

What is the value of g(3)? g(x) = x^2 + 1, so what is g(3) ?

OpenStudy (anonymous):

10 .. ? @Directrix

Directrix (directrix):

Yes.

OpenStudy (anonymous):

So I just substitute 3 in for the x .. ? @Directrix

Directrix (directrix):

f(x) = 2x + 3 So, what is f(10) if f(x) = 2*x + 3

Directrix (directrix):

@sunnyday88 So, what is f of 10 =

OpenStudy (anonymous):

f(10) =2(3)+3 = 9 .. @Directrix

Directrix (directrix):

No. You are stuck on the 3. Use it only once to find g(x) and then use what you get for g(10) to find fx).

Directrix (directrix):

f (g(3)) = f ( x2 + 1) = f (3^2 + 1) = f(10) = 2x + 3 = 2*10 + 3 = ?

Directrix (directrix):

>>> f(10) =2(3)+3 = 9 .. f(10) = 2*10 + 3 = ?

OpenStudy (anonymous):

23 .. ?

Directrix (directrix):

Yes.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Yes, 23 is correct. for this: If f(x) = 2x + 3 and g(x) = x2 + 1,find f(g(3)).

OpenStudy (anonymous):

23(10(3)) .. = 690 @Directrix

Directrix (directrix):

23 is the answer. Please follow the steps above. @sunnyday88 If there is a problem, post the answer options so that we can see what form they are in.